OpenStudy (anonymous):

what is 6(x+3)-(x-10)=18

OpenStudy (anonymous):

first distribute the 6

OpenStudy (oaktree):

6x+18+x+10=18, so 7x=-10 and x=-10/7.

OpenStudy (anonymous):

\[6x+18-(x-10)=18\] you can also distribute the -1 in front of (x-10) \[6x+18-x+10=18\]

OpenStudy (anonymous):

\[5x+28=18\] (combining constants and like terms)

OpenStudy (anonymous):

subtract 28 from both sides so that the equation is still equal \[5x+28-28=18-28\] \[5x=-10\]

OpenStudy (anonymous):

im sorry. im totally confused. i havent done this in 8 years.

OpenStudy (anonymous):

divide both sides by 5 \[\frac{5x}{5}=\frac{-10}{5}\]

OpenStudy (anonymous):

x =-2.... Odalys do you remember distribution?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

alright basically distrubution is when you have something such as \[n(a+b)\] where n, a , and b can be any variable (x,y,z,a,b,c) or constant(numbers: 1,2,3,etc)

OpenStudy (anonymous):

now in distribution you take the n and multiply it to every term in the parenthesis

OpenStudy (anonymous):